solve this equation by completing the square -4r^2+21r=r+13

Answer by stanbon(75887)

solve this equation by completing the square
-4r^2+21r=r+13
----
-4r^2 + 20r - 13 = 0
----
r = [-20 +- sqrt(400 -4*-4*-13)]/(2*-4)
----
r = [-20 +- sqrt(192)]/(-8)
----
r = [-20 +- 8sqrt(2)]/(-8)
----
r = (5/2) +- sqrt(2)
========================
Cheers,
stan H.